A rectangle has an area given by A = x^2 + 5x + 6
Find an expression for the possible length of the rectangle.
please i need help asap!! thankyou

Answer by ewatrrr(24785)


Hi
A = L*W for a rectangle
A = x^2 + 5x + 6
factoring
A = (x+3)(x+2)
Either (x+3) or (x+2) would be an expression for the possible length of this rectangle